# What Matters in Transformers? Not All Attention is Needed

### v3

@@ Abstract
Scaling Transformer-based large language models (LLMs) has demonstrated promising performance across various tasks. However, it also introduces redundant structures, posing challenges for real-world deployment. Despite some recognition of redundancy in LLMs, the variability of redundancy across different modules, such as MLP and Attention layers, is under-explored. In this work, we investigate the varying redundancy across different modules within Transformers, including Blocks, MLP, and Attention layers, using a similarity-based metric. This metric operates on the premise that redundant structures produce outputs highly similar to their inputs. Surprisingly, while attention layers are essential for transformers and distinguish them from other mainstream architectures, we found that a large proportion of attention layers exhibit excessively high similarity and can be safely pruned without degrading performance, leading to reduced memory and computation costs. Additionally, we further propose a method that jointly drops Attention and MLP layers, achieving improved performance and dropping ratios. Extensive experiments demonstrate the effectiveness of our methods, e.g., Llama-3-70B maintains comparable performance even after pruning half of the attention layers. Our findings provide valuable insights for future network architecture design. The code is released at: https://github.com/Shwai-He/LLM-Drop.
@@ 1 Introduction
Transformer-based large language models (LLMs) have significantly advanced AI research, achieving remarkable performance across various domains. Notable examples include ChatGPT OpenAI (2024) and Gemini Team (2024), which have demonstrated the potential of scaling LLMs to enhance capabilities in natural language understanding and generation. However, scaling these models introduces challenges, particularly the increase in redundant modules Frantar et al. (2023); Sun et al. (2023), which inflate deployment costs and resource demands. For instance, GPT-3, with its 175 billion parameters, requires approximately 350GB in FP16 format, far exceeding the 96GB memory capacity of the latest H100 GPUs, and making deployment on edge devices impractical. Recent studies Men et al. (2024); Gromov et al. (2024) have identified that many parameters in LLMs are redundant and can be pruned without compromising performance, highlighting the need to address redundancy within these models.
Although various previous works have been proposed to optimize LLM efficiency Frantar et al. (2023); Sun et al. (2023); He et al. (2024), these efforts often overlook the unique characteristics of transformer architectures. Specifically, transformer Vaswani et al. (2023) architectures are composed of multiple stacked blocks, each containing an MLP layer and an Attention layer, which exhibit different types and levels of redundancy. This motivates us to investigate the varying redundancy across these distinct modules.
Specifically, we propose a similarity-based metric to measure redundancy, operating on the premise that redundant modules produce outputs highly similar to their inputs. By computing the similarity between inputs and outputs, we can identify and drop redundant modules. Using this metric, we first empirically investigate Block Drop on the resulting performance. Even though some blocks exhibit relatively high similarity scores (e.g., more than 90% cosine similarity), Block Drop leads to substantial performance drops.
Given that a Transformer block comprises an MLP layer and an Attention layer Vaswani et al. (2023), we examine the effects of dropping these layers separately. Our findings reveal that while dropping MLP layers negatively impacts performance, dropping Attention layers, i.e., the core of Transformer architectures which distinguish it from other mainstream architectures (e.g., RWKV Peng et al. (2023) and Mamba Gu and Dao (2024)), can be done extensively without degrading model performance. For example, dropping 50% of the Attention layers in Llama-2-70B Touvron et al. (2023) results in comparable performance to the original model, indicating a high level of redundancy in these layers.
@@ 2 Related Works
Large Language Models. Although Large Language Models (LLMs) have demonstrated promising performance across various tasks, their deployment costs still remain a significant challenge for practical usage. Transformer Vaswani et al. (2023) models consist of multiple blocks, which include attention layers and MLP layers. Attention layers compute the contextual information between input tokens with quadratic complexity concerning the input sequence length Li et al. (2020). KV-Cache Pope et al. (2022) mitigates the computational issue but results in excessive memory costs Zhang et al. (2023). MLP layers Liu et al. (2021); Mai et al. (2022) transform each token independently, using an up-projection followed by a down-projection, and account for most of the model’s parameters. Recent works have revealed that not all blocks or layers are important Men et al. (2024); Chen et al. (2024), which urges us to reflect on the architecture redundancy within LLMs and potentially more compact architectures.
Model Compression. LLMs can be compressed to promote their efficiency in memory and computation. Quantization Frantar et al. (2023); Lin et al. (2024) and Pruning Sun et al. (2023); Frantar and Alistarh (2023) are the most widely used techniques to compress LLMs. Specifically, quantization transforms the data type into low-bit but remains potentially redundant architecture and parameters. Pruning can be categorized into unstructured pruning Kusupati et al. (2020); Sanh et al. (2020) and structured pruning Zhuang et al. (2020); Kwon et al. (2020), where the former does not apply to hardware while the latter degrades the performance significantly. Compared to pruning, our methods, i.e., Block Drop and Layer Drop, not only make the model architectures more compact but also maintain the performance. On the other hand, Block Drop and Layer Drop are orthogonal to quantization, and their integration with quantization significantly enhances the efficiency.
@@ 3.2 Block Drop
Transformer models are always stacked by multiple blocks, where each block shares the same architecture and can be seen as a subnetwork. Therefore, we first consider dropping unimportant blocks.
@@
where $\boldsymbol{X}^{l}_{B}$ and $\boldsymbol{Y}^{l}_{B}$ denote the input and output of the $l$-th block, respectively. Since the similarity scores are computed locally, we can offload irrelevant modules to save memory. Specifically, we sequentially compute the importance score of each block from shallow to deep, and then drop the blocks with the lowest scores.
@@
A transformer block consists of an Attention layer and an MLP layer, each performing different functions on the input sequence, i.e., contextual information flow between tokens and token transformation. Considering the different functionalities of Attention and MLP, we discuss them separately.
MLP Drop. As illustrated in Figure 1, an MLP layer follows LayerNorm and its outputs are a part of the overall output which involves residual connections. Given the input $\boldsymbol{X}^{l}_{M}$ of the LayerNorm before MLP at the $l$-th Block, the output $\boldsymbol{Y}^{l}_{{M}}$ can be formulated as:
@@
Attention Drop. Similarly, Attention Drop also considers the residual connection, as the outputs of attention layers contribute only partially to the overall outputs:
@@
where $\boldsymbol{X}^{l}_{{A}}$ is the inputs of the corresponding LayerNorm layers and $\boldsymbol{Y}^{l}_{{A}}$ overall outputs that involves residual connections. Like MLP Drop associates LayerNorm layers together, the importance score for Attention Drop is:
@@
Similar Block Drop, the computation of importance scores of Layer Drop is conducted layer-wisely from shallow layers to deep layers. All drop methods are conducted with a calibration dataset $\mathcal{D}$ (e.g., C4 Raffel et al. (2020)) in a one-shot manner that computes the importance scores once and then drops all redundant layers, instead of using resource-consuming multi-turn iterations. The effectiveness of One-Shot Drop has been verified in Appendix B. On the other hand, other than solely dropping Attention or MLP layers individually, we propose Joint Layer Drop that considers the combined importance scores across Attention and MLP to optimize the dropping process. The algorithm for Joint Layer Drop is presented in Algorithm 1.
@@
After removing the insignificant experts, the pruned model can be easily loaded using existing packages (e.g., Huggingface Transformers Wolf et al. (2020)) with just a change of the model configuration. More implementation details are listed in the Appendix A.
(a) Block Drop.
(b) MLP Drop.
(c) Attention Drop.
Figure 3: Visualization of Dropping order for Block Drop and Layer Drop. We visualize of the remaining layers and blocks under different dropped numbers, where yellow areas represent the retained portions and red areas indicate the dropped layers/blocks.
@@
Figure 4: Visualization of Importance Scores in Checkpoints during the Pre-training Process of MAP-Neo-7B, where light areas represent low importance scores (i.e., high similarity scores). We present both (1) Early Stage (the first 12 released checkpoints) and (2) Entire Process (checkpoints for every 500B trained tokens). We independently visualize the importance scores at the module index 0, since they are significantly higher.
@@
4 Experiments
4.1 Main Results
Comparison of Dropping Different Target Modules. In Table 1, we compare dropping different modules (i.e., Block, MLP, and Attention) with fixed numbers (e.g., 4 and 8). Since Blocks are the fundamental components of transformers, we first investigate Block Drop. Block Drop significantly degrades the performance, e.g., decreasing by 2.4% and 7.5% on the averaged performance when dropping 8 Blocks on Llama-2-13B, respectively. Given the large performance gap, Block drop is sufficient to maintain the performance.
We next delve into the removal of more fine-grained modules in Blocks, i.e., MLP and Attention Layers. Similar to Block Drop, MLP Drop also leads to a substantial performance decline, e.g., 6.3% on Llama-2-13B and 6.9% on Mistral-7B when dropping 8 layers, which is unacceptable for practical usage.
Conversely, Attention Drop demonstrates surprising effectiveness, as models with 8 dropped layers still maintain most of their performance, e.g., both Llama-3-13B and Mistral-7B maintain more than 99% of the original performance. Although attention layers are the hallmarks of transformers, dropping such layers does not degrade the performance significantly.
Efficiency Promotion of Module Dropping. We also compare the practical speedup memory usage of dropping different modules. We run models with various input and generated sequence lengths (from 512 to 4096) on a single Nvidia A100 GPU and report the average cost. While dropping blocks that contain both Attention layers and MLP layers offer the best efficiency, the significant performance drop limits practical application. Compared to MLP layers, attention layers involve fewer memory costs but have more complex operations, resulting in greater contributions to speedup after dropping (1.17×1.17$\times$1.17 × v.s. 1.09×1.09$\times$1.09 × when dropping 8 layers on Mistral-7B).
(a) Evaluation of Block Drop.
(b) Evaluation of MLP Drop.
(c) Evaluation of Attention Drop.
Attention Drop Maintains the Performance. To obtain a universal understanding of module dropping, we further conduct experiments on different a series of dropping ratios (from 0% to 100%), and the results are shown in Figure 5. Transformer models exhibit sensitivity to both Block Drop and Layer Drop, where even dropping a small number of blocks or MLP layers results in a significant performance decline. However, Attention Drop initially shows a stable trend that maintains comparable performance when dropping ratios are below 50% but subsequently decreases due to excessive layer removal.
Attention Drop Significantly Reduces the KV Cache. Given the auto-regressive nature of attention, where the output is generated token by token, the key-value (KV) cache is used to store intermediate representations of input sequences to speed up the inference process by avoiding redundant computations. Despite its role in speeding up inference, the KV cache results in significant extra memory usage Zhang et al. (2023); Singhania et al. (2024). For instance, an input with a batch size of 128 and a sequence length of 2048 for Llama-2-13B would result in about 70GB of KV cache. Our proposed Attention Drop method can safely remove 20 out of 40 attention layers, effectively reducing the KV cache by half and significantly lowering the memory requirements.
@@ 4.2 Detailed Analysis
Consistent Redundancy of Attention Layers Throughout Training Progression. To investigate how high similarity patterns are achieved, we revisit the historical checkpoints to track the dynamic changing of similarity.
Specifically, we evaluate checkpoints released by MAP-Neo-7B Zhang et al. (2024), since it released continuous checkpoints during training stages. Figure 4 presents the importance scores of Blocks and Layers at different training stages, where Attention layers demonstrate excessively low importance scores at all training stages. However, for MLP layers and Blocks, the importance scores gradually increase as training progresses, indicating that they cannot be dropped as readily as Attention layers.
@@
(a) Llama-2-13B
(b) Mistral-7B
Block Drop and Layer Drop are Data Robust. In Figure 7 and Figure 11, we note that feature similarity remains relatively stable across different modules of Mistral-7B as the sample size increases (i.e., from 4 to 1024), indicating that Block Drop and Layer Drop maintain consistency regardless of sample quantity. This confirms that using $256$ samples suffice for computing similarity, which is adopted for all our experiments. Similarly, Figure 8 and Figure 12 show that varying the calibration datasets, from the pretraining dataset (e.g., C4 Raffel et al. (2020)) to instruction tuning datasets (e.g., CodeAlpaca-20k [^1] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k, MathInstruct Yue et al. (2024) and LIMA Zhou et al. (2024)), does not significantly alter feature similarity. Despite some differences, the modules dropped remain the same. This demonstrates the resilience of Block Drop and Layer Drop to variations in calibration datasets.
Figure 7: The Impact of Sample Quantity on the Importance Scores of Attention Layers.
Figure 8: The Impact of Dataset Type on the Importance Scores of Attention Layers.
@@
As shown in Figure 6, the performance of Joint Drop is noticeably higher than either Attention Drop or MLP Drop alone. Joint Drop first only drops Attention Layers until the number of dropped layers exceeds 14 for Mistral-7B and 18 for Llama-2-13B. This is why the curves for Attention Drop and Joint Drop overlap in the left part of the graph. This process demonstrates that Attention layers are more redundant than MLP layers, which is consistent with the differing performance observed between MLP Drop and Attention Drop. Notably, Joint Drop improves the performance of Attention Drop, achieving higher dropping ratios and superior performance.
Insights for Future Network Architecture Design. Despite the success of scaling up large language models, our work provides valuable insights for scaling down models to achieve more efficient architectures. On the one hand, given the high redundancy of attention layers, future works can reduce the number of attention layers (e.g. dropping deeper attention layers), rather than keeping it consistent with that of MLP layers. On the other hand, unlike MLP layers, Attention layers demonstrate consistent redundancy as training progresses, which may be a potential bottleneck in training large language models. Replacing attention layers or developing new training techniques may further enhance the capacity of language models.

### v4

@@ Abstract
Scaling Transformer-based large language models (LLMs) has demonstrated promising performance across various tasks. However, it also introduces redundant structures, posing challenges for real-world deployment. Despite some recognition of redundancy in LLMs, the variability of structured redundancy across different modules, such as MLP and Attention layers, is under-explored. In this work, we investigate the varying redundancy across different modules within Transformers, including Blocks, MLP, and Attention layers. Surprisingly, while attention layers are essential for transformers and distinguish them from other mainstream architectures, we found that a large proportion of attention layers exhibit excessively high similarity and can be safely pruned without degrading performance, leading to reduced memory and computation costs. Additionally, we further propose a novel strategy, Joint Layer Drop, that jointly drops Attention and MLP layers, achieving improved performance and dropping ratios. Extensive experiments demonstrate the effectiveness of our methods, e.g., Llama-3-70B maintains comparable performance even after pruning half of the attention layers. Our findings provide valuable insights for future network architecture design. The code is released at: https://github.com/Shwai-He/LLM-Drop.
@@ 1 Introduction
Transformer-based large language models (LLMs) have significantly advanced AI research, achieving remarkable performance across various domains OpenAI (2024); Team (2024), which have demonstrated that scaling LLMs enhances their capabilities in natural language understanding and generation. However, scaling LLMs also introduces efficiency challenges, particularly the increase in redundant modules Frantar et al. (2023); Sun et al. (2023), which inflate deployment costs and resource demands. For instance, GPT-3, with its 175 billion parameters, requires approximately 350GB in FP16 format, far exceeding the 96GB memory capacity of the latest H100 GPUs, and making deployment on edge devices impractical. Recent studies Men et al. (2024); Gromov et al. (2024) have identified that redundant structures in LLMs that can be pruned without compromising performance, highlighting the potential of addressing the structured redundancy within these models to enhance efficiency.
Although various previous works have been proposed to promote the efficiency of LLMs Frantar et al. (2023); Sun et al. (2023); He et al. (2024), these efforts often overlook the unique characteristics of transformer architectures. Specifically, transformer Vaswani et al. (2023) architectures are composed of multiple stacked blocks, each containing an MLP layer and an Attention layer, which have different functionalities and exhibit different levels of redundancy. This motivates us to investigate the varying redundancy across these distinct modules.
Specifically, we propose a unified similarity-based metric to measure redundancy within different modules, operating on the premise that redundant modules produce outputs highly similar to their inputs. By computing the similarity between inputs and outputs, we can identify and drop redundant modules. Using this metric, we first empirically investigate Block Drop and its impact on performance. Although some blocks exhibit relatively high similarity scores (e.g., more than 90% cosine similarity), Block Drop still leads to substantial performance drops.
Given that a Transformer block comprises an MLP layer and an Attention layer Vaswani et al. (2023), we further examine the effects of dropping these layers separately. Our findings reveal that while dropping MLP layers negatively impacts performance, dropping Attention layers, i.e., the core of Transformer architectures which distinguish it from other mainstream architectures (e.g., RWKV Peng et al. (2023) and Mamba Gu and Dao (2024)), can be done extensively without degrading model performance. For example, dropping 50% of the Attention layers in Llama-2-70B Touvron et al. (2023) results in comparable performance to the original model, indicating a high level of redundancy in these layers.
@@ 2 Related Works
Transformer-based Large Language Models. Although Transformer-based Large Language Models (LLMs) have demonstrated promising performance across various tasks, their deployment costs still remain a significant challenge for practical usage Sun et al. (2023); Lin et al. (2024); Gromov et al. (2024). Transformer Vaswani et al. (2023) models consist of multiple blocks, which include Attention layers and MLP layers. Attention layers compute the contextual information between input tokens with quadratic complexity concerning the input sequence length Li et al. (2020). KV-Cache Pope et al. (2022) mitigates the computational issue but results in excessive memory costs Zhang et al. (2023). MLP layers Liu et al. (2021); Mai et al. (2022) transform each token independently, using an up-projection followed by a down-projection, and contribute most of the model parameters. Recent works have revealed that not all blocks or layers are equally important Men et al. (2024); Chen et al. (2024), which urges us to reflect on the structured redundancy within LLMs and the potential design of more compact architectures.
Model Compression. LLMs can be compressed to promote their efficiency in memory and computation. Quantization Frantar et al. (2023); Lin et al. (2024) and Pruning Sun et al. (2023); Frantar and Alistarh (2023) are the most widely used techniques to compress LLMs. Specifically, quantization transforms the data type into low-bit but remains potentially redundant architecture and parameters. Pruning can be categorized into unstructured pruning Kusupati et al. (2020); Sanh et al. (2020) and structured pruning Zhuang et al. (2020); Kwon et al. (2020). While unstructured pruning maintains better performance than structured pruning, it cannot be effectively applied to hardware, limiting its practical usage. Our methods, Block Drop and Layer Drop, focus on removing structured modules rather than fine-grained parameters, creating hardware-friendly efficient architectures while maintaining comparable performance. Additionally, Block Drop and Layer Drop are orthogonal to quantization, and their integration with quantization significantly enhances efficiency.
@@ 3.2 Block Drop
Transformer models are always stacked by multiple blocks, where each block often shares the similar architecture and can be seen as a subnetwork. Therefore, we first consider dropping unimportant blocks.
@@
where $\boldsymbol{X}^{l}_{B}$ and $\boldsymbol{Y}^{l}_{B}$ denote the input and output of the $l$-th block, respectively. Since the similarity scores are computed locally, we can offload irrelevant modules to save memory. Specifically, we compute the importance score of each block sequentially from shallow blocks to deep blocks, and then drop the blocks with the lowest scores.
@@
A transformer block consists of an Attention layer and an MLP layer, each performing different functions on the input sequence, i.e., contextual information flow between tokens and token transformation. Given the different functionalities of MLP and Attention layers, we discuss them separately.
MLP Drop. As illustrated in Figure 1, an MLP layer follows LayerNorm and its outputs are a part of the overall output which involves residual connections. Given the input $\boldsymbol{X}^{l}_{M}$ of the LayerNorm before MLP at the $l$-th Block, the overall output $\boldsymbol{Y}^{l}_{{M}}$ can be formulated as:
@@
Attention Drop. Similarly, as the outputs of attention layers contribute only partially to the overall outputs, Attention Drop also considers the residual connection:
@@
where $\boldsymbol{X}^{l}_{{A}}$ is the inputs of the corresponding LayerNorm layers and $\boldsymbol{Y}^{l}_{{A}}$ is overall outputs that involves residual connections. The importance score for Attention Drop is:
@@
Similar to Block Drop, the computation of importance scores of Layer Drop is conducted layer-wisely from shallow layers to deep layers. All drop methods are conducted with a calibration dataset $\mathcal{D}$ (e.g., C4 Raffel et al. (2020)) in a one-shot manner that computes the importance scores once and then drops all redundant layers, instead of using resource-consuming multi-turn iterations. The effectiveness of One-Shot Drop has been verified in Appendix B. On the other hand, other than solely dropping Attention or MLP layers individually, we propose Joint Layer Drop that considers the combined importance scores across Attention and MLP layers to optimize the dropping process. The algorithm for Joint Layer Drop is presented in Algorithm 1.
@@
After removing insignificant modules, the pruned model can be easily loaded using existing packages (e.g., Huggingface Transformers Wolf et al. (2020)) with just a change of the model configuration. More implementation details are listed in the Appendix A.
@@
Figure 3: Visualization of Importance Scores for MAP-Neo-7B during the Pretraining Process, where light areas represent low importance scores (i.e., high similarity scores). We present both (1) Early Stage (the first 12 released checkpoints) and (2) Entire Process (checkpoints for every 500B trained tokens). We independently visualize the importance scores at the module index 0, since they are significantly higher.
4 Experiments
4.1 Main Results
Comparison of Dropping Different Target Modules. In Table 1, we compare dropping different modules (i.e., Block, MLP, and Attention) with fixed numbers (e.g., 4 and 8). Block Drop significantly degrades the performance, e.g., decreasing by 7.5% and 11.7% on the averaged performance when dropping 8 Blocks on Llama-2-13B and Mistral-7B, respectively. Given the significant performance gap, Block Drop is insufficient to maintain the performance.
We next delve into the removal of more fine-grained modules, i.e., MLP and Attention Layers. Similar to Block Drop, MLP Drop also leads to a substantial performance decline, e.g., 6.3% on Llama-2-13B and 6.9% on Mistral-7B when dropping 8 layers.
Conversely, Attention Drop demonstrates surprising effectiveness, as models with 8 dropped layers still maintain the original performance, e.g., both Llama-3-13B and Mistral-7B maintain more than 99% of the original performance. Although attention layers are the hallmarks of transformers, dropping such layers does not degrade the performance significantly.
Efficiency Promotion of Module Dropping. We also compare the practical speedup and memory usage of dropping different modules. We run models with various input and generated sequence lengths (from 512 to 4096) on a single Nvidia A100 GPU and report the average cost. While dropping blocks that contain both Attention layers and MLP layers offer the best efficiency, the significant performance drop limits practical application. Compared to MLP layers, attention layers involve fewer memory costs but have more complex operations, resulting in greater speedup when dropped (1.17×1.17$\times$1.17 × vs. 1.09×1.09$\times$1.09 × when dropping 8 layers on Mistral-7B).
(a) Evaluation of Block Drop.
(b) Evaluation of MLP Drop.
(c) Evaluation of Attention Drop.
@@ 4.1 Main Results
(a) Block Drop.
(b) MLP Drop.
(c) Attention Drop.
Figure 5: Visualization of Dropping Order for Block Drop and Layer Drop. We visualize the remaining layers and blocks under varying dropped numbers, where yellow areas represent the retained layers/blocks and red areas indicate the dropped ones.
Attention Drop Maintains the Performance. To obtain a universal understanding of module dropping, we further conduct experiments on different a series of dropping ratios (from 0% to 100%), and the results are shown in Figure 4. Transformer models exhibit sensitivity to both Block Drop and Layer Drop, where even dropping a small number of blocks or MLP layers results in a significant performance degradation. However, Attention Drop shows a stable trend that maintains comparable performance when dropping ratios are below 50%.
Attention Drop Reduces the KV Cache Significantly. In the auto-regressive generation process, where outputs are generated token by token, the key-value (KV) cache is employed to store intermediate representations for previous tokens in attention layers, which avoids redundant computations and accelerates inference. However, the KV cache incurs substantial additional memory usage Zhang et al. (2023); Singhania et al. (2024). For instance, a batch of 128 sequences with a length of 2048 for Llama-2-13B requires approximately 70GB of KV cache. Our proposed Attention Drop method enables the removal of 20 out of 40 attention layers, effectively halving the KV cache size and significantly reducing memory requirements.
@@ 4.2 Detailed Analysis
Consistent Redundancy of Attention Layers throughout Training Progression. To investigate how the high similarity patterns are achieved, we revisit the historical checkpoints to track the dynamic changing of similarity.
Specifically, we evaluate checkpoints released by MAP-Neo-7B Zhang et al. (2024), since it has released continuous checkpoints during training stages. Figure 3 presents the importance scores of Blocks and Layers at different training stages, where Attention layers demonstrate excessively low importance scores throughout all training stages. However, the importance scores of MLP layers and Blocks gradually increase as training progresses, indicating that they cannot be dropped as readily as Attention layers.
Figure 6: Impact of Sample Quantity on the Importance Scores of Attention Layers.
Figure 7: Impact of Dataset Type on the Importance Scores of Attention Layers.
Block Drop and Layer Drop are Data Robust. In Figure 6 and Figure 11 show that the distribution of similarity scores remains stable, even when the sample size of the calibration dataset varies (from 4 to 1024). This stability indicates that Block Drop and Layer Drop maintain consistency regardless of the sample quantity, with $256$ samples being sufficient for computing similarity, as used in all our experiments.
Additionally, Figure 7 and Figure 12 demonstrate that changing the type of calibration dataset—ranging from pretraining datasets (e.g., C4 Raffel et al. (2020)) to instruction tuning datasets (e.g., CodeAlpaca-20k [^1] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k, MathInstruct Yue et al. (2024), and LIMA Zhou et al. (2024))—does not significantly impact the similarity scores distribution. Despite minor variations, the modules dropped remain nearly the same, showcasing the robustness of Block Drop and Layer Drop to changes in calibration datasets.
(a) Llama-2-13B
(b) Mistral-7B
@@
As shown in Figure 8, the performance of Joint Drop is noticeably higher than either Attention Drop or MLP Drop alone. Joint Drop first only drops Attention Layers until the number of dropped layers exceeds 14 for Mistral-7B and 18 for Llama-2-13B. This is why the curves for Attention Drop and Joint Drop overlap in the left part of the graph. Once a substantial proportion of Attention layers are removed, some MLP layers become relatively more redundant and are subsequently dropped. This process demonstrates that Attention layers are more redundant than MLP layers, which is consistent with the differing performance observed between MLP Drop and Attention Drop. Notably, Joint Drop improves the performance of Attention Drop, achieving higher dropping ratios and superior performance.
Insights for Future Network Architecture Design. Despite the successes of scaling up large language models, our work offers valuable insights into scaling down models to achieve more efficient architectures. On one hand, due to the high redundancy observed in attention layers, future works could focus on reducing the number of attention layers, particularly in deeper layers. On the other hand, On the other hand, the consistent redundancy of attention layers throughout training—potentially due to current training techniques—indicates that optimizing the training process could unlock their full potential. Additionally, exploring alternatives to attention layers may further enhance the efficiency and capacity of language models.